16. Johansen Corporation uses a predetermined overhead rate based on direct labor-hours to apply manufacturing overhead...

16.

Johansen Corporation uses a predetermined overhead rate based on direct labor-hours to apply manufacturing overhead to jobs. The Corporation has provided the following estimated costs for the next year:

Direct materials $ 6,000
Direct labor $ 20,000
Rent on factory building $ 15,000
Sales salaries $ 25,000
Depreciation on factory equipment $ 8,000
Indirect labor $ 12,000
Production supervisor's salary $ 15,000

Jameson estimates that 20,000 direct labor-hours will be worked during the year. The predetermined overhead rate per hour will be:

  • $2.50 per direct labor-hour

  • $2.79 per direct labor-hour

  • $3.00 per direct labor-hour

  • $4.00 per direct labor-hour

17. Rawe Corporation’s accounts receivable at the end of Year 2 was $341,000 and its accounts receivable at the end of Year 1 was $293,000. Sales, all on account, amounted to $1,395,000 in Year 2. The company’s average collection period for Year 2 is closest to: (Round your intermediate calculations to 2 decimal places.)

  • 89.2 days

  • 83.0 days

  • 76.7 days

  • 4.4 days

18.

Data concerning three of the activity cost pools of Salcido LLC, a legal firm, have been provided below:

Activity Cost Pool Total Cost Total Activity
Researching legal issues $ 22,580 780 research hours
Meeting with clients $ 1,303,209 7,533 meeting hours
Preparing documents $ 93,940 5,880 documents

The activity rate for the “meeting with clients” activity cost pool is closest to:

  • $76 per meeting hour

  • $173 per meeting hour

  • $1,303,209 per meeting hour

  • $97 per meeting hour

19. Which of the following would be an acceptable measure of activity for a material handling activity cost pool?

Options       Number of material moves Weight of material moved
A Yes Yes
B No Yes
C Yes No
D No No
  • Option D

  • Option C

  • Option A

  • Option B
Answer #1

16) Overhead rate = Estimated overhead/Estimated labor hour = 50000/20000 = 2.50 per labor hour

So answer is a) $2.50 per direct labor-hour

17) Average account receivable = (341000+293000/2) = 317000

Average collection period = 317000*365/1395000 = 82.9 Days

So answer is b) 83.0 days

18) Meeting with client rate = 1303209/7533 = 173

So answer is b) $173 per meeting hour

19) Which of the following would be an acceptable measure of activity for a material handling activity cost pool?

Answer is c) Option A